# BatteryML: An Open-source Platform for Machine Learning on Battery Degradation

### v4

@@ Abstract
Battery degradation remains a pivotal concern in the energy storage domain, with machine learning emerging as a potent tool to drive forward insights and solutions. However, this intersection of electrochemical science and machine learning poses complex challenges. Machine learning experts often grapple with the intricacies of battery science, while battery researchers face hurdles in adapting intricate models tailored to specific datasets. Beyond this, a cohesive standard for battery degradation modeling, inclusive of data formats and evaluative benchmarks, is conspicuously absent. Recognizing these impediments, we present BatteryML—a one-step, all-encompass, and open-source platform designed to unify data preprocessing, feature extraction, and the implementation of both traditional and state-of-the-art models. This streamlined approach promises to enhance the practicality and efficiency of research applications. BatteryML seeks to fill this void, fostering an environment where experts from diverse specializations can collaboratively contribute, thus elevating the collective understanding and advancement of battery research.
@@ 1 Introduction
Confronting this degradation complexity, reliably predicting a battery’s remaining useful life (RUL), state of health (SOH), and state of charge (SOC) becomes a herculean endeavor (Lipu et al., 2018). The significance of RUL, especially in battery management, second-hand vehicle evaluation, and more, has spurred extensive research. For instance, integrating techniques like electrochemical impedance spectroscopy with machine learning has been demonstrated to hold promise (Zhang et al., 2020; Severson et al., 2019; Attia et al., 2021; 2020). Similarly, SOH and SOC estimation has seen advances through capacity-based, Coulomb counting, impedance methods, and model-based techniques, with machine learning bringing innovative dimensions (Wang et al., 2011; Plett, 2004; Barsoukov et al., 2005; Doyle et al., 1993; He et al., 2011).
@@
[…]
@@ 3 BatteryML Platform
As depicted in Figure 1, the BatteryML pipeline comprises an organized sequence of functional modules, guiding users through the process of model creation and application. The initial step involves converting all incoming data into a consistent format. Following this, a configuration file is crafted to specify data locations, partitioning strategies, feature and label generation methods, as well as the associated model parameters. An elaborate sample of these settings is presented in Code D.1. Once configured, the comprehensive pipeline process allows end-to-end battery degradation modeling. Components integrated in this process encompass the train-test data split module, label extractor module, feature extractor module, data preprocessing module, and the model module. Detailed examples of the pipeline’s core functions, including pipeline initialization, model training, and result extraction, is available in Appendix D.1.
  • Train-test split module. This module determines the split strategy of the learning process. BatteryML allows users to randomly allocate battery cells into training and test subsets with respect to a given split proportion. Alternatively, BatteryML also provide standard data splits for popular datasets such as MATR Severson et al. (2019) and HUST Ma et al. (2022) to enable reproducible and comparable experiments. The module also offers high flexibility for custom data partitioning.
@@
  • Feature extractor module. This module implements the popular features designed by domain experts such as the cycle-difference features for battery life prediction Severson et al. (2019). This module also enables extraction of raw electric signals and conversion to tensors for training neural networks. This module is highly extensible to support custom feature design.
@@ 3.1 BatteryData: A Unified Battery Data Representation
A comprehensive outline of this BatteryData is presented in Appendix  D.2.
@@ 3.2 Feature Engineering
Both the within- and between-cycle features are represented as tensors that are compatible with modern machine learning frameworks. This allows a flexible manipulation of the feature space and a natural combination with deep learning models. We provide a detailed introduction of BatteryML feature module in Appendix D.3. Through these feature extraction methodologies, BatteryML empowers users to freely design and combine features, significantly simplifying the process of reproducing existing work and testing new features.
@@ 3.4 Model Development
Following  Attia et al. (2021), BatteryML incorporates multiple off-the-shelf baselines for battery lifetime predictions, including linear models, tree-based models and neural networks. Domain-Enhanced methods like the ’Variance’, ’Discharge’, and ’Full’ models from the  Severson et al. (2019) are implemented using handcrafted features and linear models. BatteryML also includes statistical models such as Ridge regression Hoerl & Kennard (2000), Principal Component Regression(PCR) Tipping & Bishop (1999), Partial Least Squares Regression(PLSR) Geladi & Kowalski (1986), Gaussian process Williams & Rasmussen (2006), XGBoost Chen & Guestrin (2016), Random forest  Breiman (2001) models. For high-performance needs, we offer neural network models like Multi-Layer Perceptron (MLP) Haykin (1994), Convolutional Neural Networks (CNN) Krizhevsky et al. (2012), Long Short-Term Memory networks (LSTM) Hochreiter & Schmidhuber (1997). Additionally, we introduce Transformer Vaswani et al. (2017), a groundbreaking architecture in language and vision domains  (Brown et al., 2020; Dosovitskiy et al., 2021), as a new neural baseline. These implementations utilize scikit-learn Pedregosa et al. (2011) for all statistical models, barring XGBoost, and PyTorch Paszke et al. (2019) for neural networks. We re-train each model with 10 random seeds and report averaged results to eliminate the effect of random initializaiton.
To cater to the diversity of experimental contexts, users have the liberty to tweak and customize these models as per their needs. A deeper dive into model intricacies can be found in AppendixD.5.
@@ 4.1 Data
We based our evaluation on several publicly accessible battery datasets: CALCE (Xing et al., 2013; Liu et al., 2017), HNEI (Devie et al., 2018), HUST (Ma et al., 2022), MATR (Severson et al., 2019; Hong et al., 2020), RWTH (Li et al., 2021), SNL (Preger et al., 2020), and UL_PUR (Juarez-Robles et al., 2020; 2021). These datasets encompass LFP, LCO, and NMC battery types. Further details are outlined in Table 1. The datasets differ in terms of materials, capacities, voltages, and RUL ranges. For RUL tasks, we also created combined datasets from the public sources to assess training efficacy when various battery data are combined. Notably, CRUH combines CALCE, RWTH, UL_PUR, and HNEI datasets; CRUSH merges CALCE, RWTH, UL_PUR, SNL, and HNEI datasets; and MIX incorporates all datasets used in our study. For more detailed information on the data, please refer to the AppendixA.
@@
HNEI LCO/graphite 2.8 3.0-4.3 248±15 64±17 49±28 14
RWTH NMC/graphite 1.11 3.5-3.9 658±64 60±24 46±22 48
SNL LFP/graphite 1.1 2.0-3.6 1256±1321 86±7 45±27 61
UL_PUR LCO/graphite 3.4 2.7-4.2 209±50 89±6 41±33 10
@@ 4.2 Battery Degradation Modeling
Please refer to the appendix B.1 for more detailed comparison analysis. We also provided an in-depth exploration of the impact of features and model hyperparameters in appendix C.
@@
The effectiveness of methods in SOH prediction varies across datasets. Linear models are generally effective but face challenges with the MATR cells due to variable charging strategies. Tree-based models show consistent, robust performance across datasets, establishing a strong baseline in SOH estimation. Deep learning models, however, haven’t consistently outperformed traditional methods, indicating potential areas for improvement. We provided detailed analysis in appendix B.2.
@@
LightGBM consistently surpassed other methodologies in most tasks, thereby establishing tree-based models as the current state-of-the-art in SOC prediction. Moreover, linear models continue to excel over deep learning models, highlighting the need for further research to unlock the full potential of neural networks in battery modeling. For detailed insights, please see the appendixB.3.
@@ 5 Conclusion
At the core of BatteryML is a commitment to fostering collaboration and bridging divides. As a comprehensive open-source platform, it effectively bridges the knowledge chasm between battery researchers and AI experts, streamlining data preprocessing, feature extraction, and model application, both traditional and advanced. This synthesis not only elevates battery modeling endeavors but also catalyzes a two-way exchange, i.e., empowering battery scientists to harness AI-driven tools for research and equipping AI professionals with insights to tackle intricacies specific to the battery sector.
@@ Appendix A Benchmark Dataset Construction
CALCE data source.
@@
MATR data source.
@@
HUST data source.
@@
HNEI data source.
@@
UL_PUR data source.
@@
RWTH data source.
@@
Appendix B Detailed Comparison Analysis
@@
B.1 Remaining Useful Life Prediction
@@
These models achieve significant results by fitting on low-level features, such as raw QdLinear curves. Ridge Regression Hoerl & Kennard (2000) displays consistent, albeit modest, performance across datasets, affirming a strong linear correlation between Qd(Vd) feature and battery cycle life. However, its underperformance on datasets like HUST and CRUSH, underscores the challenges in linear correlation learning. Principal Component Regression(PCR Tipping & Bishop (1999) and artial Least Squares Regression(PLSR) Geladi & Kowalski (1986), which apply principal component analysis and project cells into different subspaces before fitting a linear regression, excel in MATR1, CRUH, and SNL datasets. This suggests that the Qd(Vd) feature exhibits a linear relationship in different subspaces with the battery life. Gaussian Process Regression (GPR) Williams & Rasmussen (2006) underperforms PCR and PLSR on all datasets, indicating that the pairwise kernel functions struggle to capture the effective degradation patterns for all kinds of batteries. XGBoost Chen & Guestrin (2016) and Random Forest Breiman (2001), exhibiting strong performance on CRUSH and MIX datasets, indicate that tree-based models’ non-linear capabilities are adept at effectively modeling complex battery aging patterns.
@@
Neural network models, including Multilayer Perceptron (MLP) Haykin (1994), Convolutional Neural Network (CNN) Krizhevsky et al. (2012), Long Short-Term Memory (LSTM) Hochreiter & Schmidhuber (1997), and Transformer Vaswani et al. (2017) architectures, display significant performance variability. MLP performs well on datasets such as SNL, CLO, and CRUH, while CNN shows high sensitivity to initial conditions. LSTM, exhibiting less variability, proves its robustness and superior performance on the MATR2 dataset among the neural network methods.
@@
B.2 State of Health Estimation
@@
The SOH in this experiment is quantified as a percentage, representing the ratio of a battery’s maximum discharge capacity to its nominal capacity per cycle. For example, a battery with a nominal capacity of 2Ah and a current maximum discharge capacity of 1.8Ah equates to an SOH of $\frac{1.8}{2}\times 100=90\%$, yielding an experimental SOH range of 0 to 100.
This task leverages datasets from an array of sources, including CALCE, HNEI, HUST, MATR, RWTH, SNL, and UL_PUR. Feature extraction is comprehensive, encompassing elements such as the ratio of the current cycle’s maximum charge capacity to the nominal capacity, and the voltage of the initial cycle. We scrutinized a variety of models, including Random Forest, Linear Regression, Ridge Regression, and Neural Networks like Gated Recurrent Unit(GRU) Cho et al. (2014), MLP, and LSTM. Each model was run with 10 different seeds on specific data, so the values in each cell of the table represent the mean and standard deviation (std) of the RMSE results from multiple seeds. The benchmark results of the SOH estimation task are delineated in Table 4.
@@
B.3 State of Charge Estimation
Estimating the State of Charge (SOC) is contingent upon both the discharge capacity at a given moment and the SOH for the current cycle. Consequently, acquiring completely accurate SOC labels is as challenging as it is for SOH. In our methodology, we predict the SOC under the realistic workload of the battery. For this task, we employed datasets from CALCE, HNEI, HUST, MATR, RWTH, SNL, and UL_PUR. Feature extraction incorporates estimated current, voltage, and time information post-interpolation, alongside the charge and discharge capacity curves from preceding cycles. Labels represent the proportion of the battery’s remaining capacity at a specific point in a cycle to the current full battery capacity, multiplied by 100. The specific definition is in Formula LABEL:eq:soc_and_soh. The performance of the methods of SOC prediction task is displayed in Table 5.

### v5

@@ Abstract
Battery degradation remains a pivotal concern in the energy storage domain, with machine learning emerging as a potent tool to drive forward insights and solutions. However, this intersection of electrochemical science and machine learning poses complex challenges. Machine learning experts often grapple with the intricacies of battery science, while battery researchers face hurdles in adapting intricate models tailored to specific datasets. Beyond this, a cohesive standard for battery degradation modeling, inclusive of data formats and evaluative benchmarks, is conspicuously absent. Recognizing these impediments, We present BatteryML [^1] Project repository: https://github.com/microsoft/BatteryML - a one-step, all-encompass, and open-source platform that integrates data preprocessing, feature extraction, and the implementation of both conventional and state-of-the-art models. This streamlined approach promises to enhance the practicality and efficiency of research applications. BatteryML seeks to fill this void, fostering a collaborative platform where experts from diverse specializations can contribute, thereby accelerating collective progress in battery research.
@@ 1 Introduction
Confronting this degradation complexity, reliably predicting a battery’s remaining useful life (RUL), state of health (SOH), and state of charge (SOC) becomes a herculean endeavor (Lipu et al., 2018). The significance of RUL, especially in battery management, second-hand vehicle evaluation, and more, has spurred extensive research. For instance, integrating techniques like electrochemical impedance spectroscopy with machine learning has been demonstrated to hold promise (Zhang et al., 2020; Severson et al., 2019; Attia et al., 2021; 2020). Similarly, SOH and SOC estimation has seen advances through capacity-based, Coulomb counting, impedance methods, and model-based techniques, with machine learning bringing innovative dimensions (Wang et al., 2011; Plett, 2004; Barsoukov et al., 2005; Doyle et al., 1993; He et al., 2011b).
@@
  • Data heterogeneity. Battery data exhibits considerable heterogeneity in terms of both data format and data patterns. The output format of various battery testing systems differs with respect to the recorded fields, time granularity, file types, etc. Sometimes, severe conceptual confusion may arise due to differences in terminology conventions. For example, the capacity of the battery may be reported through the areal specific capacity, the total capacity, or even normalized capacity. Subsequent data processing further adds to the data heterogeneity. Even for the same data format, different cathode material compositions such as LiCoO2 (LCO), LiFePO4 (LFP), and LiNiMnCoO2 (NMC) lead to diverse degradation patterns.
@@ 3 BatteryML Platform
As depicted in Figure 1, the BatteryML pipeline comprises an organized sequence of functional modules, guiding users through the process of model creation and application. The initial step involves converting all incoming data into a consistent format. Following this, a configuration file is crafted to specify data locations, partitioning strategies, feature and label generation methods, as well as the associated model parameters. An elaborate sample of these settings is presented in Code E.1. Once configured, the comprehensive pipeline process allows end-to-end battery degradation modeling. Components integrated in this process encompass the train-test data split module, label extractor module, feature extractor module, data preprocessing module, and the model module. Detailed examples of the pipeline’s core functions, including pipeline initialization, model training, and result extraction, is available in Appendix E.1.
  • Train-test split module. This module determines the split strategy of the learning process. BatteryML allows users to randomly allocate battery cells into training and test subsets with respect to a given split proportion. Alternatively, BatteryML also provide standard data splits for popular datasets such as MATR (Severson et al., 2019) and HUST (Ma et al., 2022) to enable reproducible and comparable experiments. The module also offers high flexibility for custom data partitioning.
@@
  • Feature extractor module. This module implements the popular features designed by domain experts such as the cycle-difference features for battery life prediction (Severson et al., 2019). This module also enables extraction of raw electric signals and conversion to tensors for training neural networks. This module is highly extensible to support custom feature design.
@@ 3.1 BatteryData: A Unified Battery Data Representation
A comprehensive outline of this BatteryData is presented in Appendix E.2.
@@ 3.2 Feature Engineering
Both the within- and between-cycle features are represented as tensors that are compatible with modern machine learning frameworks. This allows a flexible manipulation of the feature space and a natural combination with deep learning models. We provide a detailed introduction of BatteryML feature module in Appendix E.3. Through these feature extraction methodologies, BatteryML empowers users to freely design and combine features, significantly simplifying the process of reproducing existing work and testing new features.
@@ 3.4 Model Development
Following  (Attia et al., 2021), BatteryML incorporates multiple off-the-shelf baselines for battery lifetime predictions, including linear models, tree-based models and neural networks. Domain-Enhanced methods like the ’Variance’, ’Discharge’, and ’Full’ models from the  (Severson et al., 2019) are implemented using handcrafted features and linear models. BatteryML also includes statistical models such as Ridge regression (Hoerl & Kennard, 2000), Principal Component Regression(PCR) (Tipping & Bishop, 1999), Partial Least Squares Regression(PLSR) (Geladi & Kowalski, 1986), Gaussian process (Williams & Rasmussen, 2006), XGBoost (Chen & Guestrin, 2016), Random forest  (Breiman, 2001) models. For high-performance needs, we offer neural network models like Multi-Layer Perceptron (MLP) (Haykin, 1994), Convolutional Neural Networks (CNN) (Krizhevsky et al., 2012), Long Short-Term Memory networks (LSTM) (Hochreiter & Schmidhuber, 1997). Additionally, we introduce Transformer (Vaswani et al., 2017), a groundbreaking architecture in language and vision domains (Brown et al., 2020; Dosovitskiy et al., 2021), as a new neural baseline. These implementations utilize scikit-learn (Pedregosa et al., 2011) for all statistical models, barring XGBoost, and PyTorch (Paszke et al., 2019) for neural networks. We re-train each model with 10 random seeds and report averaged results to eliminate the effect of random initializaiton.
To cater to the diversity of experimental contexts, users have the liberty to tweak and customize these models as per their needs. A deeper dive into model intricacies can be found in Appendix B.
@@ 4.1 Data
We based our evaluation on several publicly accessible battery datasets: CALCE (Xing et al., 2013; He et al., 2011a), HNEI (Devie et al., 2018), HUST (Ma et al., 2022), MATR (Severson et al., 2019; Hong et al., 2020), RWTH (Li et al., 2021), SNL (Preger et al., 2020), and UL_PUR (Juarez-Robles et al., 2020; 2021). These datasets encompass LFP, LCO, NMC, NCA and NMC_LCO battery types. Further details are outlined in Table 1. Certain datasets were excluded due to their unsuitability for tasks such as RUL estimation. The datasets differ in terms of materials, capacities, voltages, and RUL ranges. For RUL tasks, we also created combined datasets from the public sources to assess training efficacy when various battery data are combined. Notably, CRUH combines CALCE, RWTH, UL_PUR, and HNEI datasets; CRUSH merges CALCE, RWTH, UL_PUR, SNL, and HNEI datasets; and MIX incorporates all datasets used in our study. For more detailed information on the data, please refer to the Appendix A.
@@
HNEI NMC_LCO/graphite 2.8 3.0-4.3 248±15 64±17 49±28 14
RWTH NMC/carbon 1.11 3.5-3.9 658±64 60±24 46±22 48
SNL NCA,NMC,LFP/graphite 1.1 2.0-3.6 1256±1321 86±7 45±27 61
UL_PUR NCA/graphite 3.4 2.7-4.2 209±50 89±6 41±33 10
@@ 4.2 Battery Degradation Modeling
Please refer to the appendix C.1 for more detailed comparison analysis. We also provided an in-depth exploration of the impact of features and model hyperparameters in appendix D.
@@
The effectiveness of methods in SOH prediction varies across datasets. Linear models are generally effective but face challenges with the MATR cells due to variable charging strategies. Tree-based models show consistent, robust performance across datasets, establishing a strong baseline in SOH estimation. Deep learning models, however, haven’t consistently outperformed traditional methods, indicating potential areas for improvement. We provided detailed analysis in appendix C.2.
@@
LightGBM consistently surpassed other methodologies in most tasks, thereby establishing tree-based models as the current state-of-the-art in SOC prediction. Moreover, linear models continue to excel over deep learning models, highlighting the need for further research to unlock the full potential of neural networks in battery modeling. For detailed insights, please see the appendix C.3.
@@ 5 Conclusion
At the core of BatteryML is a commitment to fostering collaboration and bridging divides. As a comprehensive open-source platform, it effectively bridges the knowledge chasm between battery researchers and AI experts, streamlining data preprocessing, feature extraction, and model application, both traditional and advanced. This synthesis not only elevates battery modeling endeavors but also catalyzes a two-way exchange, i.e., empowering battery scientists to harness AI-driven tools for research and equipping AI experts with insights to tackle intricacies specific to the battery sector.
@@ Appendix A Benchmark Dataset Construction
CALCE data source (Xing et al., 2013; He et al., 2011a)
@@
MATR data source (Severson et al., 2019; Attia et al., 2020)
@@
HUST data source (Ma et al., 2022)
@@
HNEI data source (Devie et al., 2018)
@@
UL_PUR data source (Juarez-Robles et al., 2020; 2021)
@@
RWTH data source (Li et al., 2021)
@@
Appendix B Detailed Model Description
Dummy Regression.
This model uses the mean of the training samples as predictions, usually acting as a trivial baseline to indicate the error upper bound. The DummyRegressor from the scikit-learn (Pedregosa et al., 2011) toolkit is adopted in BatteryML.
Linear Regression.
Elastic Net (Zou & Hastie, 2005)
Gaussian Process (Williams & Rasmussen, 2006)
Principal Component Regression (Tipping & Bishop, 1999)
Partial Least Squares Regression (Geladi & Kowalski, 1986)
Ridge Regression (Hoerl & Kennard, 2000)
Support Vector Regression (Bishop, 2006)
XGBoost Regression (Chen & Guestrin, 2016)
XGBoostRegressor is an implementation of the gradient boosting decision trees algorithm, designed for speed and performance, that is particularly suited for regression problems, providing a robust and versatile model capable of handling a variety of data types, achieving excellent predictive performance.
LightGBM Regression (Ke et al., 2017)
LightGBMRegressor is an advanced implementation of gradient boosting decision tree algorithm, developed by Microsoft, that is designed for efficiency of compute time and memory resources. It is especially suitable for handling large datasets, excels in regression problems and offers superior predictive performance. LightGBM supports both continuous and categorical features.
Random Forest (Breiman, 2001; Geurts et al., 2006)
Multilayer Perceptron (Haykin, 1994)
Long-Short-Term Memory (Hochreiter & Schmidhuber, 1997)
Gated Recurrent Unit (Cho et al., 2014)
Gated Recurrent Unit (GRU) is a variant of the recurrent neural network (RNN), designed to adaptively capture dependencies of different time scales. Like the LSTM, GRU is also used extensively in sequence prediction problems. It features gating units that modulate the flow of information inside the unit, without having separate memory cells. They have been shown to perform comparably to LSTM with less complexity and computational cost.
Convolutional Neural Networks (Krizhevsky et al., 2012)
Transformer (Vaswani et al., 2017)
Appendix C Detailed Comparison Analysis
@@
C.1 Remaining Useful Life Prediction
@@
These models achieve significant results by fitting on low-level features, such as raw QdLinear curves. Ridge Regression displays consistent, albeit modest, performance across datasets, affirming a strong linear correlation between Qd(Vd) feature and battery cycle life. However, its underperformance on datasets like HUST and CRUSH, underscores the challenges in linear correlation learning. Principal Component Regression(PCR and artial Least Squares Regression(PLSR), which apply principal component analysis and project cells into different subspaces before fitting a linear regression, excel in MATR1, CRUH, and SNL datasets. This suggests that the Qd(Vd) feature exhibits a linear relationship in different subspaces with the battery life. Gaussian Process Regression (GPR) underperforms PCR and PLSR on all datasets, indicating that the pairwise kernel functions struggle to capture the effective degradation patterns for all kinds of batteries. XGBoost and Random Forest, exhibiting strong performance on CRUSH and MIX datasets, indicate that tree-based models’ non-linear capabilities are adept at effectively modeling complex battery aging patterns.
@@
Neural network models, including Multilayer Perceptron (MLP), Convolutional Neural Network (CNN), Long Short-Term Memory (LSTM), and Transformer architectures, display significant performance variability. MLP performs well on datasets such as SNL, CLO, and CRUH, while CNN shows high sensitivity to initial conditions. LSTM, exhibiting less variability, proves its robustness and superior performance on the MATR2 dataset among the neural network methods.
@@
C.2 State of Health Estimation
@@
The SOH in this experiment is quantified as a percentage, representing the ratio of a battery’s maximum discharge capacity to its nominal capacity per cycle following the Formula 1. For example, a battery with a nominal capacity of 2Ah and a current maximum discharge capacity of 1.8Ah equates to an SOH of $\frac{1.8}{2}\times 100=90\%$, yielding an experimental SOH range of 0 to 100.
This task leverages datasets from an array of sources, including CALCE, HNEI, HUST, MATR, RWTH, SNL, and UL_PUR. Feature extraction is comprehensive, encompassing elements such as the ratio of the current cycle’s maximum charge capacity to the nominal capacity, and the voltage of the initial cycle. We scrutinized a variety of models, including Random Forest, Linear Regression, Ridge Regression, and Neural Networks like GRU, MLP, and LSTM. Each model was run with 10 different seeds on specific data, so the values in each cell of the table represent the mean and standard deviation (std) of the RMSE results from multiple seeds. The benchmark results of the SOH estimation task are delineated in Table 4.
@@
C.3 State of Charge Estimation
Estimating the State of Charge (SOC) is contingent upon both the discharge capacity at a given moment and the SOH for the current cycle. Consequently, acquiring completely accurate SOC labels is as challenging as it is for SOH. In our methodology, we predict the SOC under the realistic workload of the battery. For this task, we employed datasets from CALCE, HNEI, HUST, MATR, RWTH, SNL, and UL_PUR. Feature extraction incorporates estimated current, voltage, and time information post-interpolation, alongside the charge and discharge capacity curves from preceding cycles. Labels represent the proportion of the battery’s remaining capacity at a specific point in a cycle to the current full battery capacity, multiplied by 100. The specific definition is in Formula 2. The performance of the methods of SOC prediction task is displayed in Table 5.